OpenStudy (liliakarina):

Distance bewtween two points is measured by the formula d= sqrt(X2-X1)^2 + (Y2-Y1)^2 I have this problem: 'What's the distance between A(3,4) and B(-3,-2) I did the operation and I get this result: 6.32 But the options are: 1.- 2 u ( u stands for units) 2.- 2 sqrt 6u 3.- 4u 4.- 6 sqrt 2u What am I doing wrong?

OpenStudy (anonymous):

the answer is one of those choice... it is expressed in exact form , not the decimal form your calculator is probably giving you...

OpenStudy (anonymous):

let's see.. what is x2-x1 = y2-y1 = answer both of these...

OpenStudy (anonymous):

hello?

OpenStudy (liliakarina):

I got 6 and 2

OpenStudy (liliakarina):

Then 6^2=36, and 2^2=4

OpenStudy (liliakarina):

36+4=40, sqrt40=6.32